kamryn921:

Triangle ABC has vertices A(−3, 1), B(−3, 4), and C(−7, 1. If ∆ABC is translated according to the rule (x, y) → (x − 4, y + 3) to form ∆A′B′C′, how is the translation described with words? (3 points) 2. Where are the vertices of ∆A′B′C′ located? Show your work or explain your steps. (4 points) 3. Triangle A′B′C′ is rotated 90° clockwise about the origin to form ∆A″B″C″. Is ∆ABC congruent to ∆A″B″C″? Give details to support your answer. (3 points)

KyledaGreat:

The answer: 1. The translation rule (x, y) → (x − 4, y + 3) means that every point in the original triangle is moved 4 units to the left and 3 units up to form the new triangle. Therefore, we can describe the translation as a "horizontal shift of 4 units to the left and a vertical shift of 3 units up." 2. To find the vertices of ∆A′B′C′, we apply the translation rule to each vertex of the original triangle: - A' = (-3, 1) → (-3 - 4, 1 + 3) = (-7, 4) - B' = (-3, 4) → (-3 - 4, 4 + 3) = (-7, 7) - C' = (-7, 1) → (-7 - 4, 1 + 3) = (-11, 4) Therefore, the vertices of ∆A′B′C′ are A'(-7, 4), B'(-7, 7), and C'(-11, 4). 3. No, ∆ABC is not congruent to ∆A″B″C″. A 90° clockwise rotation about the origin changes the orientation of the triangle, so the corresponding vertices of ∆A″B″C″ are not in the same order as those of ∆ABC. Additionally, the side lengths and angles of ∆A″B″C″ are different from those of ∆ABC. Therefore, ∆ABC is not congruent to ∆A″B″C″.